Session: Oct 2012 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Badary (Slepkova O. at Sun Oct 28 11:37:01 2012) Success:

Zelenchukskaya (Lysenkova M. at Sun Oct 28 11:49:21 2012) Success:

Medicina (Roda at Sun Oct 28 11:57:41 2012) Success:

Yebes (Daniel Valero at Sun Oct 28 15:33:15 2012) Success:
Not known problems

Urumqi (Hua Zhang at Sun Oct 28 16:14:50 2012) Success:
No known problems.

Westerbork (Geert Kuper at Sun Oct 28 21:52:39 2012) Success:
No known problems.

Torun (K. Borkowsk at Mon Oct 29 06:29:10 2012) Success:
No known problems.

Effelsberg (Uwe Bach at Mon Oct 29 12:10:48 2012) Success:

Onsala (Michael Lindqvist at Mon Oct 29 12:13:46 2012) Minor failures:
Fuse fixed (see eg066e) but we started a bit late.

Hartebeesthoek (Jonathan Quick at Tue Oct 30 09:00:12 2012) Minor failures:RECE
The RCP channel of the IF distributor has a sticky programmable attenuator causing dropouts after Tsys measurements on 2 scans at 10:35 and 11:16 UT. No other known problems.

Shanghai (Bo Xia at Thu Nov 01 13:53:14 2012) Success:
We have got the pcal signal until the last ftp scan. There was something wrong with the power supply of the pcal unit.

Jodrell_Bank (Alastair Gunn at Mon Nov 12 15:47:55 2012) Success:
No known problems.

Irbene (Vladislavs Bezrukovs at Thu Dec 13 09:57:10 2012) Minor failures:
Work only RCP channel